MILLIMETER- WAVE SPECTRUM OF CARBONYL DIAZIDE IN PURSUIT OF DIAZIRINONE

\underline{BRENT K. AMBERGER}, BRIAN J. ESSELMAN, R. CLAUDE WOODS and ROBERT J. McMA-HON, The University of Wisconsin - Madison Department of Chemistry, 1101 University Avenue, Madison, WI 53705.

Pyrolysis of carbonyl diazide (CO(N_{3})_{2}) has been shown to give diazirinone (CON_{2}).^{a} While diazirione decomposes over the course of a few hours under terrestrial conditions, there is the possibility for it to exist in space. In the pursuit of obtaining a rotational spectrum for diazirinone, we have started with the rotational spectroscopy of its immediate precursor, carbonyl diazide. Carbonyl diazide is highly explosive, and requires careful synthesis.^{b} Spectra in the range of 260-360 GHz were collected at room temperature and at −60^{∘}C. Ab initio calculations at the CCSD/cc-pVDZ level predict that the conformation where both azide groups are syn to the carbonyl is preferred. A second conformation, where one azide is syn and one is anti, is calculated to lie about 2 kcal/ mol higher in energy. Pure rotational transitions for the ground state and multiple low-lying excited vibrational states of the syn- syn conformation are readily observed and assigned.

aX. Zeng, H. Beckers, H. Willner and J. F. Stanton, Angew. Chem. Int. Ed. 50 (2011), 1720-1723

bA. M. Nolan, B. K. Amberger, B. J. Esselman, V. S. Thimmakondu, J. F. Stanton, R. C. Woods, and R. J. McMahon, Inorg.

Chem. 51 (2012), 9846-9851